Three people are selected for a committee from 4 Democrats and 3 Republicans. What is the probability the committee contains both Democrats and Republicans given that it contains at least two Democrats?



1
Expert's answer
2021-03-05T01:07:50-0500

Out of "4" Democrats and "3" Republicans a committee of "3" people can be select in"^7C_3" ways.

Therefore sample space "S" contain "^7C_3" number of elements.

Let "A" be an event such that at least two Democrats choose where committee contains both Democrats and Republicans.

Now out of "3" person at least "2" Democrats can be chosen such a way

(1) 2 Democrats "1" Republicans

(2) 3 Democrats

Therefore "2" Democrats and "1" Republicans can be choose in "(^4C_2\u00d7{^3C_1})" ways.

And "3" Democrats can be choose in "^4C_3" ways.

So number of elements in favour of "A" is "[(^4C_2\u00d7{^3C_1})+^4C_3]=22" .

"\therefore P(A)=\frac{22}{^7C_3}=\frac{22}{35}"

Let "B" be an event such that committee of both Democrats and Republicans.

This can be chosen in such a way

(1) "1" Democrat "2" Republicans

(2) "2" Democrats "1" Republican

First can be choose in "(^4C_1\u00d7{^3C_2})" ways and second can be choose in "(^4C_2\u00d7{^3C_1})" ways.

We have to find "P(B\/A)" .

Now "P(B/A)=\frac{P(B\cap A)}{P(A)}"

Here number of elements in favour of "(B\cap A)" is "(^4C_2\u00d7{^3C_1})=18" .

"\therefore P(B\cap A)=\frac{18}{^7C_3}=\frac{18}{35}" .

Hence "P(B/A)=\frac{P(B\cap A)}{P(B)}=\frac{\frac{18}{35}}{\frac{22}{35}}=\frac{18}{22}=\frac{9}{11}".

Which is the required probability.